# Understanding Reader Perception Shifts upon Disclosure of AI Authorship

### v1

@@ 3.3. Participants
We recruited 264 participants from a Japanese crowdsourcing platform  [^2] https://crowdworks.jp/. The task was described as “a survey to evaluate the quality and impression of written messages”, concealing the true purpose of the study. After the initial pre-disclosure evaluation, the study’s real objective was revealed. Participants were then given the option to withdraw with full compensation (600 JPY). Three participants withdrew, leaving 261 who gave additional consent for the post-disclosure evaluation.
Our experimental design included 6 acts of writing, 3 scenarios per act, and 11 levels of disclosed AI authorship (0% to 100% in 10% increments), resulting in 198 unique conditions. We aimed to collect at least 5 responses for each condition, so a total of $6~\text{acts}\times 3~\text{scenarios}\times 11~\text{AI levels}\times 5~\text{response samples}=990~\text{responses}$.
Each of the 261 participants evaluated four randomly selected combinations, yielding 1,044 collected responses. To ensure a balanced dataset across all conditions, we randomly down-sampled any condition with more than five responses to exactly five. This resulted in the final dataset of 990 responses used for our analysis.
@@ 3.4. Data Analysis
We used a mixed-methods approach to analyze the data, combining quantitative modeling of rating shifts with qualitative analysis of participants’ free-form explanations.
@@ 3.4.1. Quantitative Analysis
To analyze how the disclosure of AI authorship affected participants’ impressions, we fit a series of linear mixed-effects models.
The dependent variables in our models were the five metrics capturing the change in perception, calculated as the score differences (i.e., post-disclosure evaluation - pre-disclosure evaluation): TrustworthinessShift, CaringShift, CompetenceShift, LikabilityShift, and FutureShift.
@@
  • Act: The six acts of writing Convince, Interact, Reflect. Describe, Explore, and Imagine, with Describe set as the reference category, as it is arguably the most objective and least emotionally-valenced form.
@@
  • ApplyAI, DetectAI, PersuasionAI: The participants’ scores on the three AI literacy sub-scales of the MAILS questionnaire.
Our models included interaction terms between AIRatio and all other independent variables. We integrated participant-level random effects to account for individual differences in rating behavior and used the Akaike Information Criterion (AIC) for model selection.
@@ 3.4.2. Qualitative Analysis
We analyzed 990 open-ended responses using a multi-stage process that combined computational topic discovery and manual thematic coding.
First, we pre-processed the data by filtering out 61 non-substantive comments (e.g., “none,”, “no change”), leaving 929 responses for analysis. We then tokenized the text using fugashi (McCann, 2020) and a custom stopword list to filter out low-information words. We then used this cleaned data for an initial phase of inductive thematic discovery. We calculated multilingual Sentence-BERT embeddings (Reimers and Gurevych, 2020) and used BERTopic (Grootendorst, 2022) for embedding-based clustering with c-TF-IDF to identify potential topic structures. To determine the optimal number of topics k, we assessed topic coherence ($c_{v}$) for k values ranging from 3 to 30. We selected $k=14$ as it yielded the highest score.
This automated step produced a preliminary thematic structure, but manual review revealed semantic overlaps and misclassified responses. We therefore proceeded to a second phase of manual refinement. We merged and redefined the initial 14 topics to create a more robust and unambiguous codebook consisting of nine final themes (see Table 4). With this finalized codebook, we obtained an initial categorization of all 929 responses using GPT-5. Two of the authors then independently reviewed this classification, and any disagreements were resolved through discussion to reach a consensus.
Finally, to characterize each of the nine themes, we extracted representative key phrases from the original Japanese responses using 2-grams and 3-grams. For each theme, we concatenated all associated responses and identified the top n-grams based on their TF-IDF scores. These n-grams were then reviewed and combined to create meaningful descriptors for each theme.
We conducted the analysis described above using Python libraries including sentence-transformers, BERTopic, scikit-learn (KMeans), fugashi, pandas, and statsmodels.
@@ 4.1. Effect of Communicative Purpose on Reader Perception (RQ1)
We computed marginal and conditional $R^{2}$ to evaluate the explanatory power of the linear mixed-effect models. The marginal $R^{2}$ was approximately 0.20 across all models, indicating that the fixed effects explained a moderate amount of variance. The conditional $R^{2}$ was approximately 0.30, confirming that including participant-level random effects improved the model fit.
Table 3 summarizes the key coefficients. Full results are available in the Appendix (Tables 711). Figure 2 shows perception shifts by AIRatio and Act, grouped by reported AI literacy. Describe is set as the reference category of Act, based on which the coefficients for the other five types of acts are estimated. All models use AIRatio and Act as independent variables.
AIRatio showed a significant negative correlation in all the models ($p<.001$). As the stated percentage of AI involvement increased, perceptions of the author declined significantly across all five metrics.
@@
Figure 2. Act-level visualization of the regression analysis for all outcome variables. For visibility, the score of the perception shift was divided by the number of items to uniformize the range across the metrics. AIRatio shows a negative slope across all metrics. Notable findings: (1) Interact has a significantly negative intercept for TrustworthinessShift, CompetenceShift, CaringShift, and Likability. (2) Convince and Imagine have significantly positive intercepts for CompetenceShift. (3) Explore shows a significantly positive effect for CaringShift, and Imagine shows a significantly positive conditional effect with AIRatio for CaringShift.:
The communicative purpose (Act) of the text also had a significant influence, with effects varying across the different perception metrics. Interact had a significant negative effect on all dependent variables except FutureShift ($p<.001$ for all metrics except TrustworthinessShift, where $p<.01$). This suggests AI authorship is negatively viewed for social and interpersonal communication. In contrast, AI disclosure was viewed more favorably in other contexts. Convince and Imagine were positively associated with CompetenceShift ($p<.01$ for both). Explore was linked to a more positive shift in perceived caring (CaringShift, $p<.01$). The model also revealed a significant positive interaction between AIRatio and Imagine for CaringShift ($p<.05$), suggesting that the negative effect on perceived caring was mitigated as AI’s stated contribution increased in imaginative writing.
Taken together, these results indicate that readers were more tolerant of AI authorship in argumentative, creative, and exploratory writing compared to interpersonal communication.
Table 3. Coefficient matrix summarizing the results of the linear mixed-effects models predicting the shift in author perception based on the disclosed AI ratio, the act of writing, and the reader’s AI literacy.
@@
     To Convince 0.085 -1.058 1.496** 0.206
     To Explore 0.536 1.652** 0.247 1.384
     To Imagine 1.080 0.465 1.335** -0.090
     To Interact -1.999** -2.984*** -1.853*** -3.800***
     To Reflect -0.388 -0.048 0.801 -1.094
@@
     Apply AI 0.523 0.406 0.220
     Detect AI 0.473 0.335 0.361 0.854**
     AI Persuasion 0.048 0.105
@@
     AI Ratio : To Convince -0.383
     AI Ratio : To Explore 0.905
     AI Ratio : To Imagine 1.400*
     AI Ratio : To Interact -0.482
     AI Ratio : To Reflect 1.207
@@
     AI Ratio : Apply AI Literacy 0.337 0.478**
@@
     AI Ratio : AI Persuasion Literacy 0.435* 0.224**
@@ 4.2. Effect of AI Literacy on Reader Perception (RQ2)
As shown in Table 3, participants’ self-reported AI literacy levels, along with their interactions with AIRatio, were significant predictors of perception shifts. Specifically, higher ApplyAI scores were associated with a positive shift in perceived competence (CompetenceShift) and had a significant positive interaction with AIRatio on this metric ($p<.05$). DetectAI had significant positive main and interaction effects on the shift in author likability (LikabilityShift, $p<.01$ for both), as well as a significant positive interaction effect on the shift in trustworthiness (TrustworthinessShift, $p<.01$). Finally, the interaction between PersuasionAI and AIRatio had a significant positive influence on both the shifts in caring (CaringShift, $p<.05$) and desire for future interaction (FutureShift, $p<.01$). Overall, participants with higher AI literacy were more tolerant and sometimes even appreciative of AI use in writing.
@@ 4.3. Qualitative Reasons for Perception Shifts (RQ3)
In our thematic analysis of participants’ 929 open-ended responses, we identified nine themes: five explaining negative shifts and four explaining positive or neutral ones, which are summarized in Table 4.
@@ 4.3.1. Negative Themes: Why Perceptions Worsened
This theme captures a felt loss of authenticity and emotional connection. Participants reported that disclosure drained the text of its perceived warmth, even when they had initially found it compelling. One commented: Even though the writing is skillful, it doesn’t feel heartfelt” (Act: Imagine, AIRatio: 30%). Another explained: The text seemed well-written, but learning it was from an AI made it feel like it lacked a soul, so I lowered my evaluation” (Act: Convince, AIRatio: 20%). These comments highlight that for many readers, AI authorship severs the link to a human author’s perceived emotions and sincerity, undermining the text’s relational value.
@@
This theme centers on linguistic cues that participants associated with AI, such as a ”mechanical” or ”unnatural” style. As one participant noted, It felt more mechanical, which changed my impression of its warmth and friendliness” (Act: Convince, AIRatio: 100%). Another pointed out the importance of human oversight: I felt there were too many unnatural expressions. Using AI can be good, but without proper quality control, it can have the opposite effect” (Act: Reflect, AIRatio: 40%). This theme shows that beyond a general feeling of inauthenticity, specific stylistic artifacts perceived as non-human can directly trigger negative judgments.
@@
This was the largest negative theme. Here, participants questioned the author’s credibility and expertise upon learning of AI’s role. They expressed distrust in the content, as one participant noted: It seemed like AI explained most of the reasoning, so I found it hard to trust the author. It’s better to limit AI use to just making the writing clearer” (Act: Explore, AIRatio: 30%). In some cases, the AI’s perceived competence overshadowed the human’s: “I got the sense that AI handled some of the more technical parts… I was surprised by how intelligent the AI was, enough to actually change my impression of the writer” (Act: Describe, AIRatio: 50%). This theme reveals a key tension: while participants see value in using AI for information gathering, they grow skeptical of the author’s diligence and expertise when AI’s role seems too substantial, especially if they suspect a lack of fact-checking: “Using AI to obtain specialized information is very effective, but since it’s unclear whether the author has actually verified that information, it becomes very difficult to trust their judgment” (Act: Describe, AIRatio: 80%). This suggests that when AI is perceived to contribute core knowledge, it can undermine the author’s perceived competence.
@@
Finally, many participants viewed extensive AI use as a sign of laziness or a lack of personal effort from the author. Participants felt the author was outsourcing their responsibility to think and write. One commented, ”Using AI to write gives the impression of just being lazy” (Act: Convince, AIRatio: 40Others felt it diminished the author’s voice: “Since AI was used for about 30% of the writing, I felt the author’s own intention came through a bit less” (Act: Reflect, AIRatio: 30%). This judgment became more severe as the disclosed proportion of AI authorship increased. A participant who saw a 70% disclosure commented: “They’re relying way too much on AI, and there’s barely any effort to proofread what the AI wrote. I can’t help but lower my evaluation of them as a person” (Act: Describe, AIRatio: 70%). Indeed, 77 of the 103 responses in this theme corresponded to conditions where the disclosed AIRatio was 50% or higher, showing that readers penalize what they perceive as an over-reliance on automation.
@@ 4.3.2. Positive and Neutral Themes
The most common positive theme was the perception that AI was used in an appropriate, assistive, or supplementary capacity. This was particularly true when the disclosed AI contribution was low. 125 of these responses (68%) were for texts with an AIRatio under 50%. In such cases, participants felt the author remained in control, as one noted: “Since only about 10% of the text was edited by AI, and it was an explanation of how to operate industrial equipment, it felt like the AI simply identified and filled in missing parts of what a human had written. So, knowing that AI was used didn’t change my evaluation” (Act: Describe, AIRatio: 10%). AI was seen as a helpful partner rather than a replacement: “AI supported the writing nicely, which slightly improved my evaluation. It gave a friendly impression, as if the AI were helping the human” (Act: Reflect, AIRatio: 30%). Others saw a constructive partnership: “In the parts handled by AI, I got the impression that there were more technical terms. I felt like AI took charge of the technical side while the human handled the emotional side” (Act: Reflect, AIRatio: 40%).This suggests that when AI is perceived as a tool that supports rather than supplants the author, its use is seen as acceptable and even beneficial.
@@
This theme reflects an appreciation for AI’s ability to improve the quality and clarity of the writing. Participants valued AI’s contribution when it led to clearer explanations and better structure. One commented, “It felt like the AI provided clear and natural supplementary explanations that made the text easier to understand” (Act: Describe, AIRatio: 50%). Another appreciated the impact on conciseness: “It seems that the sentences became shorter after the AI’s editing… having shorter sentences makes it easier to follow” (Act: Explore, AIRatio: 70%). This appreciation extended to complex and creative tasks, where AI was seen as a valuable tool for articulating ideas more effectively: “Organizing and expressing complex subjects objectively isn’t easy, [so] if AI can help with that, it’s actually valuable” (Act: Explore, AIRatio: 60%). For a creative piece, one participant remarked, “Without the parts added by AI, the story would have felt shallow and lacking substance. With the AI’s support, it turned into a much better story” (Act: Imagine, AIRatio: 60%). This theme shows that AI’s contribution was valued when it was perceived to enhance the author’s message without supplanting their core intent.
@@
This theme was characterized by participants’ general surprise, interest, or admiration for AI’s capabilities. Comments were often brief and expressive of a ”wow” factor: “It’s unbelievably well done” (Act: Reflect, AIRatio: 90%) and “My overall evaluation didn’t really change, but I was impressed by how knowledgeable the AI seemed” (Act: Describe, AIRatio: 30%). For some, this positive reaction translated into a desire to engage with AI themselves: “This is the first time I read a poem written by AI. Next time, I’d like to try writing a poem with the help of AI” (Act: Imagine, AIRatio: 70%). Notably, participants whose responses fell into this theme tended to have higher self-reported AI literacy. Their average scores for ApplyAI, DetectAI, and PersuasionAI were 9.9 (SDSD=5.5), 2.2 (SDSD=4.2), and 2.7 (SDSD=3.9), respectively, compared to the overall sample averages of 4.6 (SDSD=7.7), -1.5 (SDSD=3.9), and -0.5 (SDSD=3.4)5. This aligns with our quantitative finding that higher AI literacy correlates with more positive reactions to AI authorship. Interestingly, the frequent expression of surprise from this more AI-literate subgroup suggests that the capabilities of current generative models continue to exceed the expectations of even relatively experienced users.
@@
This theme exclusively contains reactions to the 0% AIRatio condition, where participants were told the text was entirely human-written. While some simply stated that their evaluation did not change, many expressed increased appreciation for the author’s skill and sincerity upon this confirmation. One participant wrote, “Since the report was written entirely by a human without using AI, I felt (the author was) very intelligent, capable, and sincere, so I rated it highly” (Act: Explore, AIRatio: 0%). This suggests that explicitly labeling a text as ”human-written” can enhance an author’s standing, akin to the premium placed on products marketed as “handmade”.

### v2

@@ 3.3. Participants
We recruited 264 participants from a Japanese crowdsourcing platform  [^2] https://crowdworks.jp/. The task was initially described as “a survey to evaluate the quality and impression of written messages”, concealing the true purpose of the study. After the initial pre-disclosure evaluation, the true objective was revealed, and participants were given the option to withdraw with full compensation (600 JPY). Three participants withdrew, leaving 261 who gave additional consent for the post-disclosure evaluation.
Our experimental design included 6 acts of writing, 3 scenarios per act, and 11 levels of disclosed AI authorship (0% to 100% in 10% increments), resulting in 198 unique conditions. We aimed to collect at least 5 responses per condition, so a total of $6~\text{acts}\times 3~\text{scenarios}\times 11~\text{AI levels}\times 5~\text{response samples}=990~\text{responses}$.
Each participant evaluated four randomly selected conditions, yielding 1,044 collected responses. To ensure a balanced dataset across all conditions, we randomly down-sampled any condition with more than five responses to exactly five, resulting in a final dataset of 990 responses.
@@ 3.4. Data Analysis
We used a mixed-methods approach to analyze the data, combining quantitative modeling of rating shifts with qualitative analysis of participants’ open-ended explanations.
@@ 3.4.1. Quantitative Analysis
To analyze how the disclosure of AI authorship affected participants’ impressions, we fit a series of linear mixed-effects models. The dependent variables were the five metrics capturing the change in perception, calculated as score differences (i.e., post-disclosure evaluation - pre-disclosure evaluation): TrustworthinessShift, CaringShift, CompetenceShift, LikabilityShift, and FutureShift.
@@
  • Act: The six acts of writing–Convince, Interact, Reflect. Describe, Explore, and Imagine–with Describe set as the reference category due to its objective and emotionally neutral nature.
@@
  • ApplyAI, DetectAI, PersuasionAI: Participants’ scores on the three MAILS AI literacy sub-scales.
Our models included interaction terms between AIRatio and all other independent variables. We integrated participant-level random effects to account for individual differences in rating behavior and used the Akaike Information Criterion (AIC) and Likelihood Ratio Test (LRT) for model selection.
@@ 3.4.2. Qualitative Analysis
We analyzed the 990 open-ended responses using a multi-stage process combining computational topic discovery and manual thematic coding.
First, we pre-processed the data by filtering out 61 non-substantive comments (e.g., “none,”, “no change”), leaving 929 responses. We then tokenized the text using fugashi (McCann, 2020) and applied a custom stopword list to remove low-information words. We then used this cleaned data for an initial phase of inductive thematic discovery. We calculated multilingual Sentence-BERT embeddings (Reimers and Gurevych, 2020) and used BERTopic (Grootendorst, 2022) for embedding-based clustering with c-TF-IDF to identify potential topic structures. To determine the optimal number of topics k, we assessed topic coherence ($c_{v}$) for values ranging from 3 to 30. We selected $k=14$ which yielded the highest score.
This automated step produced a preliminary thematic structure, but manual review revealed semantic overlaps and misclassified responses. We therefore proceeded to a second phase of manual refinement. We merged and redefined the initial 14 topics to create a more robust and unambiguous codebook consisting of 9 final themes (see Table 4). With this finalized codebook, we obtained an initial categorization of all 929 responses using GPT-5. Two authors independently reviewed this classification, resolving disagreements through discussion.
Finally, to characterize each theme, we extracted representative key phrases from the original Japanese responses using 2-grams and 3-grams. For each theme, we concatenated all associated responses and identified the top n-grams based on their TF-IDF scores. These n-grams were then reviewed and combined to create meaningful descriptors for each theme.
We conducted the analysis described above using several Python libraries including sentence-transformers, BERTopic, scikit-learn (KMeans), fugashi, pandas, and statsmodels.
@@ 4.1. Effect of Communicative Purpose on Reader Perception (RQ1)
We computed marginal and conditional $R^{2}$ to evaluate the explanatory power of the linear mixed-effect models. The marginal $R^{2}$ was approximately 0.20 across all models, indicating that the fixed effects accounted for a moderate amount of variance. The conditional $R^{2}$ was approximately 0.30, confirming that the inclusion of participant-level random effects improved the model fit.
Table 3 summarizes the key coefficients, with full results provided in the Appendix (Tables 711). Figure 2 shows perception shifts by AIRatio and Act, grouped by reported AI literacy. The Describe act serves as the reference category for estimating coefficients for the other five communicative acts. All models use AIRatio and Act as independent variables.
@@
Figure 2. Act-level visualization of the regression analysis for all outcome variables. For visibility, the score of the perception shift was divided by the number of items to uniformize the range across the metrics. AIRatio shows a negative slope across all metrics. Notable findings: (1) Interact has a significantly negative intercept for TrustworthinessShift, CaringShift, and Likability, while it has a significantly positive intercept for CompetenceShift. (2) Convince and Imagine have significantly positive intercepts for CompetenceShift. (3) Explore shows a significantly positive effect for CaringShift, and Imagine shows a significantly positive conditional effect with AIRatio for CaringShift.
Act-level visualization of the regression analysis for all outcome variables. For visibility, the score of the perception shift was divided by the number of items to uniformize the range across the metrics. AIRatio shows a negative slope across all metrics. Notable findings: (1) Interact has a significantly negative intercept for TrustworthinessShift, CaringShift, and Likability, while it has a significantly positive intercept for CompetenceShift. (2) Convince and Imagine have significantly positive intercepts for CompetenceShift. (3) Explore shows a significantly positive effect for CaringShift, and Imagine shows a significantly positive conditional effect with AIRatio for CaringShift.
AIRatio exhibited a significant negative correlation across all models ($p<.001$). As the disclosed percentage of AI involvement increased, perceptions of the author declined significantly across all five metrics. The communicative purpose (Act) of the text also had a significant influence, with effects varying across the different perception metrics. Interact had a significant negative effect on TrustworthinessShift, CaringShift, and LikabilityShift ($p<.001$ for all of them). This suggests AI authorship is negatively viewed for social and interpersonal communication. In contrast, AI disclosure was viewed more favorably in terms of perceived competence and caring for certain acts. Convince, Interact, and Imagine were positively associated with CompetenceShift ($p<.01$ for Convince and Imagine, and $p<.001$ for Interact). Explore was linked to a more positive shift in perceived caring (CaringShift, $p<.01$). The model also revealed a significant positive interaction between AIRatio and Imagine for CaringShift ($p<.05$), suggesting that the negative effect on perceived caring was mitigated as the stated AI contribution increased in imaginative writing.
Collectively, these results indicate that readers were more tolerant of AI authorship in argumentative, creative, and exploratory writing compared to interpersonal communication.
Table 3. Coefficient matrix summarizing the results of the linear mixed-effects models predicting the shift in author perception based on the disclosed AIRatio, the act of writing, and the reader’s AI literacy. Significant effects are highlighted with color, orange for positive and blue for negative effects.
@@
     To Convince 0.085 -1.075 1.496** 0.206
     To Explore 0.536 1.612** 0.247 1.384
     To Imagine 1.080 0.410 1.335** -0.090
     To Interact -1.999** -3.050*** 1.853*** -3.800***
     To Reflect -0.388 -0.144 0.801 -1.094
@@
     Apply AI Literacy 0.523 0.220
     Detect AI Literacy 0.473 0.511 0.361 0.854**
     AI Persuasion Literacy -0.001 0.105
@@
     AIRatio : To Convince -0.390
     AIRatio : To Explore 0.892
     AIRatio : To Imagine 1.372*
     AIRatio : To Interact -0.508
     AIRatio : To Reflect 1.098
@@
     AIRatio : Apply AI Literacy 0.478**
@@
     AIRatio : AI Persuasion Literacy 0.467* 0.224**
@@ 4.2. Effect of AI Literacy on Reader Perception (RQ2)
As shown in Table 3, participants’ self-reported AI literacy levels, along with their interactions with AIRatio, were significant predictors of perception shifts. Specifically, higher ApplyAI scores were associated with a positive shift in perceived competence (CompetenceShift) and had a significant positive interaction with AIRatio on this metric ($p<.05$). DetectAI demonstrated significant positive main and interaction effects on the shift in author likability (LikabilityShift, $p<.01$ for both), as well as a significant positive interaction effect on the shift in trustworthiness (TrustworthinessShift, $p<.01$). Finally, the interaction between PersuasionAI and AIRatio had a significant positive influence on both the shifts in caring (CaringShift, $p<.05$) and desire for future interaction (FutureShift, $p<.01$). Overall, participants with higher AI literacy were more tolerant, and sometimes even appreciative, of AI use in writing.
@@ 4.3. Qualitative Reasons for Perception Shifts (RQ3)
In our thematic analysis of the 929 open-ended responses, we identified nine themes: five explaining negative shifts and four explaining positive or neutral ones. These themes are summarized in Table 4.
@@ 4.3.1. Negative Themes: Why Perceptions Worsened
This theme captures a perceived loss of authenticity and emotional connection. Participants reported that disclosure drained the text of its warmth, even when they had initially found it compelling. One commented: Even though the writing is skillful, it doesn’t feel heartfelt” (Act: Imagine, AIRatio: 30%). Another explained: The text seemed well-written, but learning it was from an AI made it feel like it lacked a soul, so I lowered my evaluation” (Act: Convince, AIRatio: 20%). These comments highlight that for many readers, AI authorship severs the link to a human author’s perceived emotions and sincerity, undermining the relational value of the text.
@@
This theme centers on linguistic cues that participants associated with AI, such as a “mechanical” or “unnatural” style. As one participant noted, It felt more mechanical, which changed my impression of its warmth and friendliness” (Act: Convince, AIRatio: 100%). Another pointed out the importance of human oversight: I felt there were too many unnatural expressions. Using AI can be good, but without proper quality control, it can have the opposite effect” (Act: Reflect, AIRatio: 40%). This theme shows that beyond a general feeling of inauthenticity, specific stylistic artifacts perceived as non-human can directly trigger negative judgments.
@@
This was the most prevalent negative theme. Participants questioned the author’s credibility and expertise upon learning of AI’s role, expressing distrust in the content. As one participant noted: It seemed like AI explained most of the reasoning, so I found it hard to trust the author. It’s better to limit AI use to just making the writing clearer” (Act: Explore, AIRatio: 30%). In some cases, the perceived competence of the AI overshadowed that of the human: “I got the sense that AI handled some of the more technical parts… I was surprised by how intelligent the AI was, enough to actually change my impression of the writer” (Act: Describe, AIRatio: 50%). This theme reveals a key tension: while participants see value in using AI for information gathering, they grow skeptical of the author’s diligence and expertise when the role of AI appears too substantial, particularly if they suspect a lack of fact-checking. One participant remarked: “Using AI to obtain specialized information is very effective, but since it’s unclear whether the author has actually verified that information, it becomes very difficult to trust their judgment” (Act: Describe, AIRatio: 80%). This suggests that when AI is perceived to contribute core knowledge, it can undermine the author’s perceived competence.
@@
Finally, many participants viewed extensive AI use as a sign of laziness or a lack of personal effort from the author. Participants felt the author was outsourcing their responsibility to think and write. One commented, ”Using AI to write gives the impression of just being lazy” (Act: Convince, AIRatio: 40%). Others felt it diminished the author’s voice: “Since AI was used for about 30% of the writing, I felt the author’s own intention came through a bit less” (Act: Reflect, AIRatio: 30%). This judgment became more severe as the disclosed proportion of AI authorship increased. A participant who saw a 70% disclosure commented: “They’re relying way too much on AI, and there’s barely any effort to proofread what the AI wrote. I can’t help but lower my evaluation of them as a person” (Act: Describe, AIRatio: 70%). Notably, 77 of the 103 responses in this theme corresponded to conditions where the disclosed AIRatio was 50% or higher, showing that readers penalize what they perceive as an over-reliance on automation.
@@ 4.3.2. Positive and Neutral Themes
The most frequent positive theme was the perception that AI was used in an appropriate, assistive, or supplementary capacity. This was particularly true when the disclosed AI contribution was low. 125 of these responses (68%) were for texts with an AIRatio under 50%. In such cases, participants felt the author remained in control, as one noted: “Since only about 10% of the text was edited by AI, and it was an explanation of how to operate industrial equipment, it felt like the AI simply identified and filled in missing parts of what a human had written. So, knowing that AI was used didn’t change my evaluation” (Act: Describe, AIRatio: 10%). AI was seen as a helpful partner rather than a replacement: “AI supported the writing nicely, which slightly improved my evaluation. It gave a friendly impression, as if the AI were helping the human” (Act: Reflect, AIRatio: 30%). Others saw a constructive partnership: “In the parts handled by AI, I got the impression that there were more technical terms. I felt like AI took charge of the technical side while the human handled the emotional side” (Act: Reflect, AIRatio: 40%). This suggests that when AI is perceived as a tool that supports rather than supplants the author, its use is deemed acceptable and even beneficial.
@@
This theme reflects an appreciation for the ability of AI to improve the quality and clarity of the writing. Participants valued the contribution of AI when it led to clearer explanations and better structure. One commented: “It felt like the AI provided clear and natural supplementary explanations that made the text easier to understand” (Act: Describe, AIRatio: 50%). Another appreciated the impact on conciseness: “It seems that the sentences became shorter after the AI’s editing… Having shorter sentences makes it easier to follow” (Act: Explore, AIRatio: 70%). This appreciation extended to complex and creative tasks, where AI was seen as a valuable tool for articulating ideas more effectively: “Organizing and expressing complex subjects objectively isn’t easy, [so] if AI can help with that, it’s actually valuable” (Act: Explore, AIRatio: 60%). For a creative piece, one participant remarked, “Without the parts added by AI, the story would have felt shallow and lacking substance. With the AI’s support, it turned into a much better story” (Act: Imagine, AIRatio: 60%). This theme shows that AI’s contribution was valued when it was perceived to enhance the author’s message without replacing their core intent.
@@
This theme was characterized by participants’ general surprise, interest, or admiration for AI’s capabilities. Comments were often brief and expressive of a “wow” factor: “It’s unbelievably well done” (Act: Reflect, AIRatio: 90%) and “My overall evaluation didn’t really change, but I was impressed by how knowledgeable the AI seemed” (Act: Describe, AIRatio: 30%). For some, this positive reaction translated into a desire to engage with AI themselves: “This is the first time I read a poem written by AI. Next time, I’d like to try writing a poem with the help of AI” (Act: Imagine, AIRatio: 70%). Participants in this category tended to have higher self-reported AI literacy. Their average scores for ApplyAI, DetectAI, and PersuasionAI were 9.9 (SDSD=5.5), 2.2 (SDSD=4.2), and 2.7 (SDSD=3.9), respectively, compared to the overall sample averages of 4.6 (SDSD=7.7), -1.5 (SDSD=3.9), and -0.5 (SDSD=3.4). This aligns with our quantitative finding that higher AI literacy correlates with more positive reactions to AI authorship. Interestingly, the frequent expressions of surprise from this more AI-literate subgroup suggest that the capabilities of current generative models continue to exceed the expectations of even relatively experienced users.
@@
This theme exclusively comprises reactions to the 0% AIRatio condition, where participants were told the text was entirely human-written. While some simply stated that their evaluation did not change, many expressed increased appreciation for the author’s skill and sincerity upon this confirmation. One participant wrote, “Since the report was written entirely by a human without using AI, I felt (the author was) very intelligent, capable, and sincere, so I rated it highly” (Act: Explore, AIRatio: 0%). This suggests that explicitly labeling a text as ”human-written” can enhance an author’s standing, akin to the premium placed on products marketed as “handmade”.